How do I set up this problem?
Mr. Fox invested a sum of money at 6%. He invested a second sum, $2000.00 less than this sum, at 5%. He invested third sum, which was $3000.00 less than the first sum, at 3%. His total annual income was $370.00. Find the amount he invested t each rate.

Answer by mananth(16949)

percent Principal $
3.00 x-3000
6.00 x
5.00 x-2000
Interest = 370
3*(x-3000)+6x+5(x-2000)= 370*100
The percentages have been multiplied by 100 for easier calculation
3x-9000+6x+5x-10000 =37000
14x=19000+37000
14x= 56000
/14
x= 4000
$ 4000 at6%
$2000 at 5%
$1000 at 3%